OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-7,-2) and (-1,-4) The equation of the line in slope-intercept form is y=?

OpenStudy (anonymous):

find the slope using the formula: get -1/3 using y-y1=m(x-x1) we have y=-1/3x-13/3

OpenStudy (anonymous):

did i do something?

OpenStudy (anonymous):

not getting the right answer unless I am inputting it wrong.